# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v7

@@ 3.2 Numerical Experiments
We conduct experiments on a dataset $S=\{1,\dots,n\}$ with $n=500$, distributed across nodes of a communication graph. Our evaluation metric is the normalized absolute error between estimated and true ranks, i.e., for node $k$ at iteration $t$, the error is defined as $\ell_{k}(t)=|R_{k}(t)-r_{k}|/n$. While more sophisticated ranking metrics such as Kendall $\tau$ exist, they are not relevant in our context, as we focus on individual rank estimation accuracy. We first examine the impact of the constants appearing in our theoretical bounds (see Theorem 1), the one related to graph connectivity and the one reflecting rank centrality. To this end, we consider three graph topologies: the complete graph ($c=4.01\times 10^{-3}$), in which every node is directly connected to all others, yielding maximal connectivity; the two-dimensional grid ($c=1.65\times 10^{-5}$), where each node connects to its four immediate neighbors; and the Watts–Strogatz network ($c=3.31\times 10^{-4}$), a randomized graph with average degree $k=4$ and rewiring probability $p=0.2$, offering intermediate connectivity between the complete and grid graphs. Then, we compare GoRank with the two other ranking algorithms Baseline and Baseline++. Figure (a) was generated on a Watts-Strogatz graph, averaged over $1e3$ trials. Figure (b) and (c) show the convergence of ranking algorithms with mean and standard deviation computed from $100$ trials: figure (b) shows the convergence of GoRank on different graph topologies; figure (c) compares the convergence of the ranking algorithms on a 2D Grid graph. All experiments were run on a single CPU with 32 GB of memory for $8e4$ iterations, with a total execution time of approximately two hours. The code for our experiments will be publicly available. [^1] The code is available in the Supplementary Material.
@@ 4.1 Algorithm – Convergence Analysis
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm˜2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation˜2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $0<c_{2}<1$ and is given by $c_{2}=c/2$ Shah et al. [2009], Boyd et al. [2006]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@ 4.2 Robustness Analysis - Breakdown Points
Consider a dataset $S$ of size $n\geq 1$ and $T_{n}=T_{n}(S)$ a real-valued statistic based on it. The breakdown point for the statistic $T_{n}(S)$ is defined as $\varepsilon^{*}=p_{\infty}/n$, where $p_{\infty}=\min\{p\in\mathbb{N^{*}}:\,\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|=\infty\}$, the supremum being taken over all corrupted datasets $S^{\prime}_{p}$ obtained by replacing $p$ samples in $S$ by arbitrary samples. The breakdown point is a popular notion of robustness Huber and Ronchetti [2011], corresponding here to the fraction of samples that need to be corrupted to make the statistic $T_{n}$ arbitrarily large (i.e., "break down"). For example, the breakdown point of the $\alpha$-trimmed mean is given by $\lfloor\alpha n\rfloor/n$, which is approximately $\alpha$. We consider here a generalization of this notion, namely the $\tau$-breakdown point by replacing $p_{\infty}$ with $p_{\tau}=\min\{p\in\mathbb{N^{*}}\enspace\text{, }\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|\geq\tau\}$, where $\tau>0$ is a threshold parameter. Since our algorithm does not compute the exact $\alpha$-trimmed mean, we focus instead on determining the $\tau$-breakdown point $\varepsilon^{*}_{k}(t)$ of the partial $\alpha$-trimmed mean at iteration $t$ for each node $k$, i.e., the estimate $Z_{k}(t)$. Given that this quantity was previously shown to converge to the $\alpha$-trimmed mean, we expect that $\varepsilon^{*}_{k}(t)\leq\alpha$. The following theorem provides framing bounds for the breakdown point of the estimates of GoTrim when paired with GoRank.
@@
$\frac{1}{n}\max\left(\left\lfloor\lfloor\alpha n\rfloor+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{k}(t)\leq\frac{\lfloor\alpha n\rfloor}{n}\enspace,$
@@ 4.3 Numerical Experiments
The experimental setup is identical to that of the previous section, with the key difference being the introduction of corrupted data. Specifically, the dataset $S$ is contaminated by replacing a fraction $\varepsilon=0.1$ of the values with outliers. We consider two types of corruption, each affecting $\lfloor\varepsilon n\rfloor$ randomly selected data points: (a) scaling, where a value $x$ is changed to $sx$, and (b) shifting, where $x$ becomes $x+s$. While this is a relatively simple form of corruption, it is sufficient to break down the classical mean. We measure performance using the absolute error between the estimated and true trimmed mean. For node $k$ at iteration $t$, the error is given by $\ell(t)=(1/n)\sum_{k}|Z_{k}(t)-\bar{x}_{\alpha}|$. Experiments (a) and (b) are run on dataset $S$, corrupted with scaling $s=10$, using a Watts-Strogatz and a 2D grid graph, respectively. Experiment (c) uses the Basel Luftklima dataset, corrupted with shift $s=100$. This dataset includes temperature measurements from $n=105$ sensors across Basel. A graph with connectivity $c=4.7\times 10^{-4}$ is constructed by connecting sensors within 1 km of each other. The code and dataset for our experiments will be publicly available. [^2] The code and Basel Luktklima dataset are available in the Supplementary Material.
@@ 5 Conclusion and Discussion
Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in Appendix H. The theoretical analysis in the asynchronous setting will be carried out in an extension to this work.
@@
GoTrim naturally extends to the decentralized estimation of rank-based statistics. This includes rank statistics [Hájek and Šidák, 1999], which are key tools in data analysis—particularly for robust hypothesis testing—as well as L-statistics (such as the Winsorized mean).
@@ Acknowledgments
This research was supported by the PEPR IA Foundry grant. The authors thank the program for its funding and support.
@@ Appendix A Gossip Algorithms for Ranking
Here, we detail two gossip algorithms for ranking, which can be considered natural competitors to GoRank. The algorithm presented first was proposed in Chiuso et al. [2011] and, to our knowledge, is the only decentralized ranking algorithm documented in the literature. The algorithm presented next can be seen as a variant of the latter, incorporating the more efficient communication scheme of GoRank.
@@
Chiuso et al. propose a gossip algorithm for distributed ranking in a general (connected) network Chiuso et al. [2011]. They demonstrate that this algorithm solves the ranking problem almost surely in finite time. However, they do not provide any convergence rate or non-asymptotic convergence results. The algorithm is outlined in Algorithm˜3 and proceeds as follows. At each time step $t$, an edge $(i,j)$ is selected, and the algorithm operates in three phases:
@@ A.2 Baseline++ - Our Improved Variant Proposal
The algorithm selects an edge $(i,j)$ at each step, and the corresponding nodes check if the auxiliary ranks are consistent with their auxiliary observations. If the ordering is inconsistent, the nodes swap their auxiliary ranks. Then, each node updates its local rank if the ordering of its local observation is inconsistent with the auxiliary observation. In contrast, the algorithm proposed by Chiuso et al. only updates the local estimates when the wandering estimate returns to its originating node. This design can significantly slow down convergence in graphs with low connectivity and long return times.
@@ B.1 Properties of Gossip Matrices
The proofs of (a), (b), (d), and (e) are omitted for brevity. For more details, we refer the reader to Boyd et al. [2006], Colin et al. [2015]. The proof of (c) follows from the fact that
@@
The proof of (f) can be found in Colin et al. [2015] (see Lemmas 1, 2 and 3). ∎
@@ B.2 Convergence Analysis of Gossip for Averaging
Algorithm 5 Gossip algorithm for estimating the standard average Boyd et al. [2006]
@@
The proof is kept brief, as this is a standard result. For more details, we refer the reader to Boyd et al. [2006], Colin et al. [2015]. To prove this lemma, we will primarily use points (e) and (f) from ˜Lemma.
@@
The original proof can be found in Boyd et al. [2006].
@@ C.2 Proof of Theorem 2
$\mathbb{E}\left[\left|R_{k}(t)-r_{k}\right|\right]\leq\sqrt{\frac{3}{ct}}\cdot\sigma_{k}\enspace.$
@@
$R^{\prime}_{k}(t)-r_{k}=\frac{1}{t}\sum_{s=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}^{\top}_{1}(s:)\mathbf{e}_{k}\enspace,$
@@
$\big{(}R^{\prime}_{k}(t)-\bar{h}_{k}\big{)}^{2}=\frac{1}{t^{2}}\sum_{s=0}^{t-1}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\enspace,$
@@
$+\frac{1}{t^{2}}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\enspace.$
@@
$v(u)^{2}=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})\tilde{\mathbf{h}}_{k}\enspace.$
@@
$v(s)v(u)=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:}s+1)\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
@@
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\mathbb{E}\left[\tilde{\mathbf{W}}_{1}(u{:}s+1)\right]\cdot\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
@@
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]\leq\lambda_{2}(1)^{u-s}\cdot\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\enspace.$
@@
$\mathbb{E}[v(s)v(u)]\leq\lambda_{2}(1)^{u-s}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$
@@
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{t^{2}}\sum_{s<u}\lambda_{2}(1)^{u-s}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$
@@
$\sum_{s<u}\lambda_{2}(1)^{u-s}\leq\sum_{u=0}^{t-1}\sum_{d=1}^{u}\lambda_{2}(1)^{d}\leq\frac{t}{1-\lambda_{2}(1)}\enspace.$
@@
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{ct}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\leq\frac{3}{ct}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\enspace.$

### v8

@@ 3.2 Numerical Experiments
We conduct experiments on a dataset $S=\{1,\dots,n\}$ with $n=500$, distributed across nodes of a communication graph. Our evaluation metric is the normalized absolute error between estimated and true ranks, i.e., for node $k$ at iteration $t$, the error is defined as $\ell_{k}(t)=|R_{k}(t)-r_{k}|/n$. While more sophisticated ranking metrics such as Kendall $\tau$ exist, they are not relevant in our context, as we focus on individual rank estimation accuracy. We first examine the impact of the constants appearing in our theoretical bounds (see Theorem 1), the one related to graph connectivity and the one reflecting rank centrality. To this end, we consider three graph topologies: the complete graph ($c=4.01\times 10^{-3}$), in which every node is directly connected to all others, yielding maximal connectivity; the two-dimensional grid ($c=1.65\times 10^{-5}$), where each node connects to its four immediate neighbors; and the Watts–Strogatz network ($c=3.31\times 10^{-4}$), a randomized graph with average degree $k=4$ and rewiring probability $p=0.2$, offering intermediate connectivity between the complete and grid graphs. Then, we compare GoRank with the two other ranking algorithms Baseline and Baseline++. Figure (a) was generated on a Watts-Strogatz graph, averaged over $1e3$ trials. Figure (b) and (c) show the convergence of ranking algorithms with mean and standard deviation computed from $100$ trials: figure (b) shows the convergence of GoRank on different graph topologies; figure (c) compares the convergence of the ranking algorithms on a 2D Grid graph. All experiments were run on a single CPU with 32 GB of memory for $8e4$ iterations, with a total execution time of approximately two hours. The code for our experiments is publicly available. [^1] The code is available at github.com/anna-vanelst/robust-gossip.
@@ 4.1 Algorithm – Convergence Analysis
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm˜2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation˜2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $0<c_{2}<1$ and is given by $c_{2}=c/2$ [shah2009gossip, boyd2006randomized]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@ 4.2 Robustness Analysis - Breakdown Points
Consider a dataset $S$ of size $n\geq 1$ and $T_{n}=T_{n}(S)$ a real-valued statistic based on it. The breakdown point for the statistic $T_{n}(S)$ is defined as $\varepsilon^{*}=p_{\infty}/n$, where $p_{\infty}=\min\{p\in\mathbb{N^{*}}:\,\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|=\infty\}$, the supremum being taken over all corrupted datasets $S^{\prime}_{p}$ obtained by replacing $p$ samples in $S$ by arbitrary samples. The breakdown point is a popular notion of robustness [huber2011robust], corresponding here to the fraction of samples that need to be corrupted to make the statistic $T_{n}$ arbitrarily large (i.e., "break down"). For example, the breakdown point of the $\alpha$-trimmed mean is given by $\lfloor\alpha n\rfloor/n$, which is approximately $\alpha$. We consider here a generalization of this notion, namely the $\tau$-breakdown point by replacing $p_{\infty}$ with $p_{\tau}=\min\{p\in\mathbb{N^{*}}\kern 5.0pt\text{, }\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|\geq\tau\}$, where $\tau>0$ is a threshold parameter. Since our algorithm does not compute the exact $\alpha$-trimmed mean, we focus instead on determining the $\tau$-breakdown point $\varepsilon^{*}_{k}(t)$ of the partial $\alpha$-trimmed mean at iteration $t$ for each node $k$, i.e., the estimate $Z_{k}(t)$. Given that this quantity was previously shown to converge to the $\alpha$-trimmed mean, we expect that $\varepsilon^{*}_{k}(t)\leq\alpha$. The following theorem provides framing bounds for the breakdown point of the estimates of GoTrim when paired with GoRank.
@@
$\frac{1}{n}\max\left(\left\lfloor\lfloor\alpha n\rfloor+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{k}(t)\leq\frac{\lfloor\alpha n\rfloor}{n}\kern 5.0pt,$
@@ 4.3 Numerical Experiments
The experimental setup is identical to that of the previous section, with the key difference being the introduction of corrupted data. Specifically, the dataset $S$ is contaminated by replacing a fraction $\varepsilon=0.1$ of the values with outliers. We consider two types of corruption, each affecting $\lfloor\varepsilon n\rfloor$ randomly selected data points: (a) scaling, where a value $x$ is changed to $sx$, and (b) shifting, where $x$ becomes $x+s$. While this is a relatively simple form of corruption, it is sufficient to break down the classical mean. We measure performance using the absolute error between the estimated and true trimmed mean. For node $k$ at iteration $t$, the error is given by $\ell(t)=(1/n)\sum_{k}|Z_{k}(t)-\bar{x}_{\alpha}|$. Experiments (a) and (b) are run on dataset $S$, corrupted with scaling $s=10$, using a Watts-Strogatz and a 2D grid graph, respectively. Experiment (c) uses the Basel Luftklima dataset, corrupted with shift $s=100$. This dataset includes temperature measurements from $n=105$ sensors across Basel. A graph with connectivity $c=4.7\times 10^{-4}$ is constructed by connecting sensors within 1 km of each other. The code and dataset for our experiments is publicly available. [^2] The code and Basel Luktklima dataset are available at github.com/anna-vanelst/robust-gossip.
@@ 5 Conclusion and Discussion
Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in Appendix H. The theoretical analysis in the asynchronous setting is carried out in an extension to this work [van2025asynchronous].
@@
GoTrim naturally extends to the decentralized estimation of rank-based statistics. This includes rank statistics [HAJEK1999], which are key tools in data analysis—particularly for robust hypothesis testing—as well as L-statistics (such as the Winsorized mean).
@@ Acknowledgments
This research was supported by the PEPR IA Foundry and Hi!Paris ANR Cluster IA France 2030 grants. The authors thank the program for its funding and support.
Outline of the Supplementary Material
The Supplementary Material is organized as follows. Section A introduces two alternative gossip algorithms for distributed ranking: Baseline and Baseline++. In Section B, we present auxiliary results related to gossip matrices and the convergence analysis of the standard gossip algorithm for averaging. The detailed convergence analysis of GoRank, including the proofs of Theorems 1 and 2, is provided in Section C. Section D presents the convergence proofs for GoTrim, specifically Lemma 1 and Theorem 3. In Section E, we provide the robustness analysis of GoTrim (i.e., the proof of Theorem 4). The convergence analysis of ClippedGossip is covered in Section F. Section G includes additional experiments and implementation details. Section H introduces Asynchronous GoRank, along with corresponding experimental results. Section I details experiments on larger networks. Section J describes how the current framework can be extended to include network disruptions. Section K provides experiments on sparse networks. Finally, Section L offers further discussion and outlines future work.
@@ Appendix A Gossip Algorithms for Ranking
Here, we detail two gossip algorithms for ranking, which can be considered natural competitors to GoRank. The algorithm presented first was proposed in [chiuso2011gossip] and, to our knowledge, is the only decentralized ranking algorithm documented in the literature. The algorithm presented next can be seen as a variant of the latter, incorporating the more efficient communication scheme of GoRank.
@@
chiuso2011gossip propose a gossip algorithm for distributed ranking in a general (connected) network [chiuso2011gossip]. They demonstrate that this algorithm solves the ranking problem almost surely in finite time. However, they do not provide any convergence rate or non-asymptotic convergence results. The algorithm is outlined in Algorithm˜3 and proceeds as follows. At each time step $t$, an edge $(i,j)$ is selected, and the algorithm operates in three phases:
@@ A.2 Baseline++ - Our Improved Variant Proposal
The algorithm selects an edge $(i,j)$ at each step, and the corresponding nodes check if the auxiliary ranks are consistent with their auxiliary observations. If the ordering is inconsistent, the nodes swap their auxiliary ranks. Then, each node updates its local rank if the ordering of its local observation is inconsistent with the auxiliary observation. In contrast, the algorithm proposed by chiuso2011gossip only updates the local estimates when the wandering estimate returns to its originating node. This design can significantly slow down convergence in graphs with low connectivity and long return times.
@@ B.1 Properties of Gossip Matrices
The proofs of (a), (b), (d), and (e) are omitted for brevity. For more details, we refer the reader to [boyd2006randomized, colin2015extending]. The proof of (c) follows from the fact that
@@
The proof of (f) can be found in [colin2015extending] (see Lemmas 1, 2 and 3). ∎
@@ B.2 Convergence Analysis of Gossip for Averaging
Algorithm 5 Gossip algorithm for estimating the standard average [boyd2006randomized]
@@
The proof is kept brief, as this is a standard result. For more details, we refer the reader to [boyd2006randomized, colin2015extending]. To prove this lemma, we will primarily use points (e) and (f) from ˜Lemma.
@@
The original proof can be found in [boyd2006randomized].
@@ C.2 Proof of Theorem 2
$\mathbb{E}\left[\left|R_{k}(t)-r_{k}\right|\right]\leq\sqrt{\frac{3}{ct}}\cdot\sigma_{k}\kern 5.0pt.$
@@
$R^{\prime}_{k}(t)-r_{k}=\frac{1}{t}\sum_{s=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}^{\top}_{1}(s:)\mathbf{e}_{k}\kern 5.0pt,$
@@
$\big(R^{\prime}_{k}(t)-\bar{h}_{k}\big)^{2}=\frac{1}{t^{2}}\sum_{s=0}^{t-1}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\kern 5.0pt,$
@@
$+\frac{1}{t^{2}}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\kern 5.0pt.$
@@
$v(u)^{2}=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:})\tilde{\mathbf{h}}_{k}\kern 5.0pt.$
@@
$v(s)v(u)=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\tilde{\mathbf{W}}_{1}(u{:}s+1)\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\kern 5.0pt.$
@@
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]=\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\mathbb{E}\left[\tilde{\mathbf{W}}_{1}(u{:}s+1)\right]\cdot\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\kern 5.0pt.$
@@
$\mathbb{E}[v(s)v(u)\mid\mathbf{W}_{1}(s{:})]\leq\lambda_{2}(1)^{u-s}\cdot\tilde{\mathbf{h}}_{k}^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\tilde{\mathbf{W}}_{1}(s{:})\tilde{\mathbf{h}}_{k}\kern 5.0pt.$
@@
$\mathbb{E}[v(s)v(u)]\leq\lambda_{2}(1)^{u-s}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\kern 5.0pt.$
@@
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{t^{2}}\sum_{s<u}\lambda_{2}(1)^{u-s}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\kern 5.0pt.$
@@
$\sum_{s<u}\lambda_{2}(1)^{u-s}\leq\sum_{u=0}^{t-1}\sum_{d=1}^{u}\lambda_{2}(1)^{d}\leq\frac{t}{1-\lambda_{2}(1)}\kern 5.0pt.$
@@
$\mathbb{E}\left[\left(R^{\prime}_{k}(t)-\bar{h}_{k}\right)^{2}\right]\leq\left(\frac{1}{t}+\frac{2}{ct}\right)\|\tilde{\mathbf{h}}_{k}\|^{2}\leq\frac{3}{ct}\cdot\|\tilde{\mathbf{h}}_{k}\|^{2}\kern 5.0pt.$